1.    The diameter of a hydrogen atom is 1.0E-10 m (i.e. 1.0E-10 m/atom). How many hydrogen atoms could fit along a 1.0 cm line?

(number = length/(unit length))

 

 

2.    Earth travels with a speed of 30 km/s around the sun. What is the circumference of Earth’s orbit (which takes one year to travel)?

(distance = speed * time)

 

3.    A light-year is the distance light travels in one year, traveling with a speed of 3.0E8 m/s. How far is that in meters? (distance = speed * time)

 

 

4.    The wavelength of an ultrasound wave is 3.0E-4 m (i.e. 3.0E-4 m/wavelength). How many wavelengths would fit in a layer of skin with thickness 2.0 mm?

(number = length/(unit length))



1
Expert's answer
2021-10-03T13:24:12-0400

1)


"N=\frac{0.01}{10^{-10}}=10^{8}"

2)


"L=30(365\cdot24\cdot3600)=9.5\cdot10^{8}\ km"

3)


"L=(3\cdot10^{8})(365\cdot24\cdot3600)=9.5\cdot10^{15}\ m"

4)


"N=\frac{0.002}{3\cdot10^{-4}}=6.7"